Use the following information to answer this question. Bayside, Inc. 2017 Income Statement ($ in thousands) Net sales Cost of goods sold Depreciation Earnings before interest and taxes Interest paid Taxable income Taxes Net income $6,920 4,390 370 $2,160 36 $2,124 743 $1,381 Bayside, Inc 2016 and 2017 Balance Sheets ($ in thousands) 2016 2017 2017 $1,615 $1,990 640 3,300 3,110 910 1,160 2016 Cash Accounts rec. Inventory Total Net fixed assets Total assets $125 $ 230 Accounts payable 1,050 1,7402,100 Common stock 890 Long-term debt 840 $2,915 $3,220 Retained earnings 3,750 3,680 $6,665 $6,900 Total liab. & equity $6,665 $6,900 What is the equity multiplier for 2017?
Answer #1

Equity Multiplier = Average Total assets / Average Shareholders Equity

Average total assets = (6665+6900)/2 = 6872.5

Average Shareholders Equity = [ (3300+910) + (3110+1160) ] /2 = 4240

Equity Multiplier = 6872.5 / 4240 = 1.62087